Calculate ionization potential and first excitation potential of the hydrogen

atom taking β„Ž = 6.62 x 10βˆ’34𝐽𝑠, 𝑒 = 1.6 x 10βˆ’19

, π‘š = 9.1 x 10βˆ’31π‘˜π‘”.


1
Expert's answer
2021-01-24T14:23:26-0500

The ionization potential of the hydrogen atom is the energy required to remove one electron at the ground state from the atom. We can use Rydberg equation:


"\frac1\lambda=R\bigg(\frac{1}{n_1^2}-\frac{1}{n_\infty^2}\bigg)=\frac{R}{n_1^2},\\\space\\
E_0=h\nu=h\frac c\lambda=hRc=2.18\cdot10^{-18}\text{ J, or 13.6 eV.}"

The first excitation potential of the hydrogen atom is


"V=\frac{3}{4}E_0=\frac34\cdot13.6=10.2\text{ eV}."